a train leaves Slaton traveling east at 80 km/h. an hour later anothertrain leaves Slaton on a parallel track at 120 km/h. how far from Slaton will the trains meet?

Answer by ankor@dixie-net.com(22740)

a train leaves Slaton traveling east at 80 km/h.
an hour later another train leaves Slaton on a parallel track at 120 km/h.
how far from Slaton will the trains meet?
:
First find the time it takes them to meet
Let t = travel time of the 1st train
then
(t-1) = travel time of the 2nd train
:
When they meet, they will have traveled the same distance,
Write a distance equation, dist = speed * time
:
120(t-) = 80t
120t - 120 = 80t
120t - 80t = 120
40t = 120
t = 3 hrs
:
Find the distance with the 1st train: 80*3 = 240 km from Slayton
Check the distance with the 2nd train: 120*2 = 240 km